### v1

# When the Correct Model Fails: The Optimality of Stackelberg Equilibria with Follower Intention Updates

Abstract
We study a two-player dynamic Stackelberg game between a leader and a follower. Classical formulations of the Stackelberg equilibrium (SE) assume that the follower’s best response (BR) mapping is known to the leader. However, this is not always true in practice. In those cases the leader needs to simultaneously infer this BR function while fulfilling an internal objective. We study a setting in which the leader selects a control strategy that optimizes an objective given an initial belief about the follower’s best response. This belief is updated during the finite decision horizon, prompting the leader to reoptimize its control. We characterize the optimality guarantees of the SE solutions under this belief update for both open loop (OL) and feedback (FB) information structures. In particular, we show that it is possible that assuming an incorrect follower BR map obtains a lower cost over the game horizon than knowing the true BR. We support these claims with numerical examples in a linear quadratic (LQ) Stackelberg game.
I Introduction
Autonomous systems have become increasingly prevalent in society over the past years, whether that is in the form of self-driving cars, delivery robots, or humanoids replacing household workers. As they become commonly deployed in environments shared by humans, vehicles, and other robots, autonomous systems’ ability to interact effectively with others become more critical to their overall performance. When the autonomous system does not know the nature of the interaction, it must be capable of learning the interaction model and optimizing the interaction strategy simultaneously. Active research areas such as intent estimation and inverse learning [1, 2, 3] implicitly assume that knowing the correct interaction model leads to better interaction outcomes. However, we re-examine this question:
does knowing the correct interaction model lead to the most optimal interaction strategy?
If the autonomous systems do not interact, i.e., if at most one autonomous system bases its strategy on the decisions of others, then that system can effectively treat the others as dynamic obstacles and perform obstacle avoidance by estimating their dynamics online. In this case, our research question becomes trivial: knowing the true model of the interaction leads better interaction outcomes [4]. However, autonomous systems inherently interact if they observe and avoid each other. Not modeling this interaction leads to incorrect outcome predictions and unsafe outcomes.
To address the posed question under autonomous interactions, this work models the interaction as a two player non-cooperative Stackelberg game [5, 6] where the leader estimates the follower’s BR and optimizes its own Stackelberg strategy under the current best BR guess simultaneously. We consider both OL and FB information structure with linear time invariant (LTI) dynamics and a finite and discrete time horizon. We focus on the problem from a strategy optimization perspective, and assume that the estimated follower’s BR comes from an oracle that may update the BR estimate during the control horizon.
Contributions. For OL Stackelberg equilibria, we prove a sufficient condition for when optimizing with respect to a correct BR leads to the optimal leader strategy (which implies that correct BR does not always lead to optimal leader strategy), and demonstrate its connection to the time inconsistency property of Stackelberg equilibria. For FB Stackelberg equilibria, we show that not all solutions imply that a correct BR leads to the optimal leader strategy despite its time consistency. We also prove that updating the leader’s strategies during the control horizon does not necessarily lead to better outcomes, even if we update with respect to a more accurate BR estimation. Finally, we demonstrate via numerical simulations of an LQ Stackelberg game with Bayesian updates of BR estimation that these theoretical results in fact frequently occur and lead to sub-optimal behavior when not accounted for.
This work opens up other questions of broader interest such as whether the metric by which we evaluate these problems should be the cumulative cost, and whether accurately learning the intentions of other autonomous systems is always worth the effort.
II Related Research
First introduced in [5] to model oligopolistic markets, the Stackelberg game and its dynamic extension [7, 8, 9] are used extensively in control theory to model lane-merging scenarios [10], pollution control [11], and power-grid optimization [12]. An important consideration in Stackelberg games is the information structure; common structures include open loop [13], feedback [14], and closed loop [15]. Both open loop and feedback Stackelberg equilibria have been analytically solved for the discrete-time linear quadratic (LQ) game [14], while [16] provides general optimality conditions for joint strategies. The closed loop Stackelberg equilibrium is more difficult to analytically derive in general. For LQ games, the two-step memory closed loop solutions are derived in [17, 15] based on a team-optimal strategy.
Although Stackelberg games assume that the leader knows the follower’s best response, this is not always true in practice. Estimating the opponents’ intentions is the focus of inverse learning [2] in dynamic game settings: [18] employs a Bayesian perspective to estimate opponent intentions, [19] uses inverse reinforcement learning [1] to model the interaction between autonomous cars and humans, and [20] provides a method to infer opponent objectives and compute a Nash equilibrium under noisy and partial observations. In particular, [18] derives an adaptive solution for finding Nash equilibria under partial observations while this paper explicitly addresses the optimality conditions of known Stackelberg equilibrium under such adaptive solutions. Finally, [21] studies a similar problem from an online learning perspective.
III The Stackelberg game under unknown follower intentions
Notation We use ${\mathbb{R}}^{n}$ to denote a real-valued vector space of dimension $n$, ${\mathbb{R}}^{n\times m}$ to denote a real-valued matrix with dimensions $n\times m$, and $\mathbb{N}$ to denote the set of all natural numbers. The discrete number set $\{0,\ldots,T\}$ is denoted as $[T]$ and the discrete number set $\{t_{0},\ldots,t_{f}\}$ is denoted as $t_{0},t_{f}]$. Additionally, we denote the state trajectory as $\mathbf{x}=(x_{0},\ldots,x_{T})\in{\mathbb{R}}^{n(T+1)}$, the control trajectory as $\mathbf{u}=(u_{0},\ldots,u_{T-1})\in{\mathbb{R}}^{mT}$, and the strategy trajectory as $\bm{\pi}=(\pi_{0},\ldots,\pi_{T-1})\in{\mathbb{R}}^{(m\times n)T}$.
III-A Stackelberg dynamic game
We consider two players, a leader $L$ and a follower $F$ over a finite and discrete time horizon $[T]$. The players jointly follow deterministic, LTI system dynamics given by
${x}_{t+1}$ $=A{x}_{t}+B^{L}u^{L}_{t}+B^{F}u^{F}_{t},\quad\forall\ t\in[T-1],$ (1)
where ${x}_{t}\in{\mathbb{R}}^{n}$ is the shared time-dependent system state. For player $i\in\{L,F\}$, $B^{i}\in{\mathbb{R}}^{n\times m_{i}},$ is its control matrix, and $u^{i}_{t}\in{\mathbb{R}}^{m_{i}}$ is its time-varying control. From initial state $x_{0}\in{\mathbb{R}}^{n}$ and under control sequences $(\mathbf{u}^{L},\mathbf{u}^{F})$, we denote the feasible state trajectory from (1) as
$\mathbf{x}=Hx_{0}+G^{L}\mathbf{u}^{L}+G^{F}\mathbf{u}^{F}\in{\mathbb{R}}^{n(T+1)},$ (2)
where the matrices $H\in{\mathbb{R}}^{n(T+1)\times n}$, $G^{i}\in{\mathbb{R}}^{n(T+1)\times m_{i}T}$ for $i\in\{L,F\}$ are defined in (44). For a time horizon $[c,d]$ with horizon length $\hat{T}=d-c+1$, we use $H_{c:d}\in{\mathbb{R}}^{n\hat{T}\times n}$, $G^{i}_{c:d}\in{\mathbb{R}}^{n\hat{T}\times m_{i}(\hat{T}-1)}$ to denote the state trajectory in horizon $[c,d]$ from state $x_{c}$ as
$x_{c:d}=H_{c:d}x_{c}+G^{L}_{c:d}{u}_{c:d-1}^{L}+G^{F}{u}_{c:d-1}^{F}\in{\mathbb{R}}^{n\hat{T}}$. (3)
Information Structure. Players select strategies to generate control based on the available information at each time step $t$, which we denote by $\mathcal{I}^{i}_{t}$ for $i\in\{L,F\}$. We consider two information structures:
  1. 1.
    open loop information $\mathcal{I}^{i}_{t}=\{x_{0}\},\forall t\in[T-1]$;
  2. 2.
    feedback information $\mathcal{I}^{i}_{t}=\{x_{t}\},\ \forall t\in[T-1]$.
Each information structure leads to a distinct strategy set [6]. We denote player $i$’s strategy set by $\Pi^{i}_{t}$. The strategy sets corresponding to the two information structures above are explicitly given by
$\Pi^{i}_{t}=\begin{cases}\mathcal{U}^{i}_{t}=\{u^{i}_{t}\in{\mathbb{R}}^{m_{i}}\},&\mathcal{I}^{i}_{t}=\{x_{0}\}\\ \mathcal{K}^{i}_{t}=\{K^{i}_{t}\in{\mathbb{R}}^{m_{i}\times n}\},&\mathcal{I}^{i}_{t}=\{x_{t}\}\end{cases},i\in\{L,F\}.$ (4)
Under FB information structure, the control chosen is $u^{i}_{t}~=~-K^{i}_{t}x_{t}$ for $i\in\{L,F\}$. We use $\Pi^{i}_{0:T-1}$ to denote player $i$’s strategy domain over the finite horizon [T1][T-1], and $\bm{\pi}^{i}\in\Pi^{i}_{0:T-1}$ to denote a specific strategy for player $i$. When the information structure context $\mathcal{I}^{i}$ is clear, we use $\bm{\pi}^{i}$ to denote $\bm{\pi}^{i}(\mathcal{I}^{i})$.
Player Objectives. At time step $t\in[T-1]$, each player $i\in\{L,F\}$ has a state and control-dependent cost function $J^{i}_{t}:{\mathbb{R}}^{n\times m_{L}\times m_{F}}\mapsto{\mathbb{R}}$. At time step $T$, players have a state-dependent cost function $J^{i}_{T}:{\mathbb{R}}^{n}\mapsto{\mathbb{R}}$. Each player’s total cost over time horizon $[T]$ is given by
$J^{i}_{0:T}\big(\mathbf{x},\mathbf{u}^{L},\mathbf{u}^{F}\big)=J^{i}_{T}\big(x_{T}\big)\\ +\sum_{t=0}^{T-1}J^{i}_{t}\big(x_{t},u^{L}_{t},u^{F}_{t}\big).$ (5)
We make strict convexity assumptions about $J^{L}_{0:T}$ and $J^{F}_{0:T}$.
Assumption 1.
The follower’s objective $J^{F}_{0:T}$ is continuously differentiable and strictly convex in $\big(\mathbf{x},\mathbf{u}^{F}\big)$ for all $\mathbf{u}^{L}\in{\mathbb{R}}^{m_{L}T}$. The leader’s objective $J^{L}_{0:T}$ is continuously differentiable and strictly convex in $\big(\mathbf{x},\mathbf{u}^{L},\mathbf{u}^{F}\big)$.
Best response (BR). Under a leader’s strategy $\bm{\pi}^{L}$, the follower seeks to minimize its objective $J^{F}_{0:T}$ (5) using its strategy $\bm{\pi}^{F}$. This problem is formulated as
$\min_{\bm{\pi}^{F}\in\Pi^{F}_{0:T-1}}$ $\quad J^{F}_{0:T}\big(\mathbf{x},\bm{\pi}^{L}(\mathcal{I}^{L}),\bm{\pi}^{F}(\mathcal{I}^{F})\big)$ (6)
s.t. $\mathbf{x}=Hx_{0}+G^{L}\bm{\pi}^{L}+G^{F}\bm{\pi}^{F}.$
Under Assumption 1, the follower has a unique best response to any leader strategy $\bm{\pi}^{L}$ [6, Sec. 7.2] over the set of injective mappings from the leader’s strategy space to the follower’s, which we denote as
$\mathcal{F}=\{b\ |\ b:\Pi_{0:T-1}^{L}\mapsto\Pi_{0:T-1}^{F}\},$ (7)
and the follower’s BR is a mapping $b^{\star}\in\mathcal{F}$ such that $b^{\star}(\bm{\pi}^{L})$ is the solution to (6) for any $\bm{\pi}^{L}\in\Pi_{0:T-1}^{L}$. Next, we define the Stackelberg dynamic game.
Definition 1 (Stackelberg Dynamic Game).
In a two player Stackelberg dynamic game with leader $L$ and follower $F$, the leader seeks a strategy $\bm{\pi}^{L}$ that solves
$\min_{\bm{\pi}^{L}\in\Pi^{L}_{0:T-1}}$ $J^{L}_{0:T}\big(\mathbf{x},\bm{\pi}^{L}(\mathcal{I}^{L}),b^{\star}(\bm{\pi}^{L})\big)$ (8)
s.t. $\mathbf{x}=Hx_{0}+G^{L}\bm{\pi}^{L}+G^{F}b^{\star}(\bm{\pi}^{L}),$
while the follower seeks a strategy $\bm{\pi}^{F}$ that solves (6). In particular, the leader always announces $\bm{\pi}^{L}$ first and the follower responds with $\bm{\pi}^{F}$.
Joint strategies that solve (8) under different information structure leads to different mathematical guarantees. We discuss each specific guarantee in later Sections IV-A and IV-B, respectively.
III-B Stackelberg game under follower BR updates
Under a known follower BR (6), we can find the SE for OL and FB information structures with guarantees [6]. To understand how simultaneous follower BR updates can affect these optimality concepts, we define a Stackelberg game with BR update in this section.
Assumption 2 (Follower’s BR).
Over the time horizon $[T]$, the follower has a time-invariant BR given by
$b^{\star}\in\mathcal{F}.$ (9)
The leader has two beliefs about the follower’s BR: at $t=0$, it is $b^{1}\in\mathcal{F}$, and at an update time $\tau\in(0,T]$, it is $b^{2}\in\mathcal{F}$.
The specific method used to retrieve the follower’s beliefs is less important here. However, many methods from estimation and machine learning can predict the follower’s BR under different observation inputs [1, 22, 2, 3].
Figure 1: Dynamics of Stackelberg game with BR update.
Stackelberg game with BR update. Under Assumption 2, consider a Stackelberg game in which the leader does not know the follower’s true belief $b^{\star}$, but the leader first solves the $[T]$-horizon Stackelberg game (8) under belief $b^{1}$ to execute strategy $\bm{\pi}^{L1}\in\Pi^{L}_{0:T-1}$, the follower solves its $[T]$-horizon problem (6) to execute strategy $b^{\star}(\bm{\pi}^{L1})$, and the system dynamics evolve under (3) for horizon $0,\tau]$ as
$x^{1}_{0:\tau}=H_{0:\tau}x_{0}+G^{L}_{0:\tau}\pi^{L1}_{0:\tau-1}+G^{F}_{0:\tau}\pi^{F1}_{0:\tau-1}.$ (10)
At the update time $\tau$, the leader solves an new optimal control problem with updated BR belief $b^{2}\in\mathcal{F}$ and shorter time horizon $\tau,T]$, given as
$\min_{\pi^{L}_{\tau:T-1}\in\Pi^{L}_{\tau:T-1}}J^{L}_{\tau:T}\left(x_{\tau:T},\pi^{L}_{\tau:T-1},b^{2}(\pi^{L}_{\tau:T-1})\right)$ (11)
$\text{s.t. }\quad x_{\tau:T}=H_{\tau:T}x_{\tau}+G^{L}_{\tau:T}\pi^{L}_{\tau:T-1}+G^{F}_{\tau:T}b^{2}(\pi^{L}_{\tau:T-1}).$
The leader computes and executes a new strategy under the new belief $b^{2}$, which we denote as $\hat{\pi}^{L2}_{\tau:T-1}$. Thus, the leader’s entire strategy over the horizon $[T]$ under BR update is
$\hat{\bm{\pi}}^{L}=\left[\pi^{L1}_{0:\tau-1},\hat{\pi}^{L2}_{\tau:T-1}\right].$ (12)
At $t=0$, the follower solves its $[T]$-horizon problem (6). When the leader updates its strategy to $\hat{\pi}^{L2}_{\tau:T-1}$ at $t=\tau$, the follower also update its best response to $b^{\star}(\hat{\pi}^{L2}_{\tau:T-1})$. Thus, the follower’s entire strategy over horizon $[T]$ is
$\hat{\bm{\pi}}^{F}=[\pi^{F1}_{0:\tau-1},b^{\star}(\hat{\pi}^{L2}_{\tau:T-1})],$ (13)
where $\pi^{F1}_{0:\tau-1}$ consists of the first $\tau$ elements of $b^{\star}(\bm{\pi}^{L1})$. Given the realized strategies $(\hat{\bm{\pi}}^{L},\hat{\bm{\pi}}^{F})$, the state trajectory in the Stackelberg game with BR update is given by
$\hat{\mathbf{x}}=[x^{1}_{0:\tau-1},x^{2}_{\tau:T}]\in{\mathbb{R}}^{n(T+1)},$ (14)
where $x^{1}_{0:\tau}$ is defined in (10) and $x^{2}_{\tau:T}$ is given by $x^{2}_{\tau:T}=H_{\tau:T}x^{1}_{\tau}+G^{L}_{\tau:T}\hat{\pi}^{L2}_{\tau:T-1}+G^{F}_{\tau:T}b^{\star}(\bm{\pi}^{L2}_{\tau:T-1})$ such that $x^{1}_{\tau}=x^{2}_{\tau}$.
To evaluate the optimality guarantees of the Stackelberg game with BR update, we compare the leader’s cost under the true BR $b^{\star}$ to the alternative beliefs $(b^{1},b^{2})\in\mathcal{F}$. To simplify the discussion, we break down the leader’s cost (5) into the components $J^{L}_{0:\tau-1}(b^{1})$ and $J^{L}_{{\tau:T}}(b^{2})$, defined as
$J^{L}_{0:\tau-1}(b^{1})$ $=\sum_{t=0}^{\tau-1}J^{i}_{t}\big(x^{1}_{t},\pi^{L1}_{t},\pi^{F1}_{t}\big),$ (15)
$J^{L}_{{\tau:T}}(b^{2})$ $=J_{T}^{i}(x^{2}_{T})+\sum_{t=\tau}^{T-1}J^{L}_{t}\big({x}^{2}_{t},\hat{{\pi}}_{t}^{L2},b^{\star}(\hat{{\pi}}_{t}^{L2})\big),$ (16)
such that $J^{L}_{0:T}(\hat{\mathbf{x}},\hat{\bm{\pi}}^{L},\hat{\bm{\pi}}^{F})=J^{L}_{0:\tau-1}(b^{1})+J^{L}_{{\tau:T}}(b^{2})$.
Having formulated the problem setting, we now state the central question addressed in this work.
Problem 1 (Stackelberg game under BR belief update).
We consider a Stackelberg game in which the leader updates its belief about the follower’s BR at time $\tau$ and the follower always utilizes BR $b^{\star}\in\mathcal{F}$. Does the true BR belief always lead to the lowest cost? I.e., is it true that for all beliefs
$J^{L}_{0:\tau-1}(b^{\star})+J^{L}_{{\tau:T}}(b^{\star})\leq J^{L}_{0:\tau-1}(b^{1})+J^{L}_{{\tau:T}}(b^{2}),\ \forall\ b^{1},b^{2}\in\mathcal{F}.$ (17)
We investigate Problem 1 for both the OL and FB information structures (4).
IV SE under BR Updates
This section addresses Problem 1 under OL and FB information structures (4). We first define the optimality conditions of both SE solutions, followed by a discussion on whether the SE solutions with BR updates and under the true BR belief satisfy (17).
IV-A Open loop Stackelberg equilibrium (OLSE) under BR updates
We first evaluate the OLSE guarantee [6, Sec.7.2] under BR updates. When the leader solves the Stackelberg game (8) against a follower whose BR is $b^{\star}\in\mathcal{F}$, the leader selects the control sequence $\mathbf{u}^{L\star}$ that satisfies
$\mathbf{u}^{L\star}$ $\in\mathop{\rm argmin}_{\mathbf{u}^{L}\in\mathcal{U}^{L}_{0:T-1}}J^{L}_{0:T}\left(\mathbf{x},\mathbf{u}^{L},b^{\star}\big(\mathbf{u}^{L}\big)\right)$ (18)
s.t. $\mathbf{x}=Hx_{0}+G^{L}\mathbf{u}^{L}+G^{F}b^{\star}\big(\mathbf{u}^{L}\big),$
for $x_{0}\in{\mathbb{R}}^{n}$. The optimality condition of the OLSE implies that the leader’s OL control $\mathbf{u}^{L\star}$ is optimal for the given initial state $x_{0}$ and with respect to all alternative control sequences $\mathbf{u}^{L}\in{\mathbb{R}}^{m_{L}T}$ announced at $t=0$ and played over the entire horizon $[T]$.
An important property of game equilibria (in particular SE) is the so-called time-inconsistency, which we now define.
Definition 2 (SE Time-inconsistency [23]).
Consider the SE $\left(\bm{\pi}^{L\star},\bm{\pi}^{F\star}\right)$ that solves (6) and (8) from initial state $x_{0}\in{\mathbb{R}}^{n}$ with information structures $(\mathcal{I}^{L},\mathcal{I}^{F})$. If at times step $t\in[T]$ and state $x_{t}\in{\mathbb{R}}^{n}$ under dynamics (2), we re-solve (6) and (8) for the remaining horizon and obtain SE $\left(\hat{\pi}^{L}_{t:T-1},\hat{\pi}^{F}_{t:T-1}\right)$. We say that the SE $\left(\bm{\pi}^{L\star},\bm{\pi}^{F\star}\right)$ is time-inconsistent if
$\pi^{L\star}_{t:T-1}\neq\hat{\pi}^{L}_{t:T-1}.$ (19)
If the SE is not time-inconsistent, then it is time-consistent.
From [24, 11, 23], the OLSE (18) is time-inconsistent.
We next consider whether $\hat{\mathbf{u}}^{L\star}$ is optimal against all other control sequences with BR updates $\hat{\mathbf{u}}^{L}$ (12). Since the OLSE is time-inconsistent, that means $\hat{\mathbf{u}}^{L\star}\neq\mathbf{u}^{L\star}$ when $\tau<T$. Thus, the optimality of (18) is not satisfied by $\mathbf{\hat{u}}^{L\star}$. We now explore what the optimality guarantees of $\mathbf{\hat{u}}^{L\star}$ are.
Under the true BR belief, at $t=0$ the leader announces $\mathbf{u}^{L\star}$ (18) for the horizon $[T]$. Then, the follower best responds to the $[T]$-horizon control $\mathbf{u}^{F\star}=b^{\star}(\mathbf{u}^{L\star})$. In the following proposition, we compare this leader strategy to all other strategies announced for a horizon $[T]$ but only played over the horizon $0,\tau)$.
Proposition 1.
Under BR update (12) with update time $\tau~<~T$ and OL information structure (4), if the $[T]$-horizon SE $\mathbf{u}^{L\star}$ (18) does not minimize the leader’s cost over the interval $0,\tau)$ (15) and initial state $x_{0}\in{\mathbb{R}}^{n}$, then there exists $b^{1}\in\mathcal{F},\ b^{1}\neq b^{\star}$ such that
$J^{L}_{0:\tau-1}(b^{1})<J^{L}_{0:\tau-1}(b^{\star}).$ (20)
Proof.
Since $\mathbf{u}^{L\star}$ does not minimize the leader’s cost over horizon $0,\tau)$, there exists a strategy announced over $[T]$ $\mathbf{v}^{L}\in{\mathbb{R}}^{m_{L}T}$ such that it minimizes the cost (15) and $\mathbf{v}^{L}~\neq~\mathbf{u}^{L\star}$. Let the follower’s BR to this control be $\mathbf{v}^{F}=b^{\star}(\mathbf{v}^{L})$.
Now, consider a belief $b^{1}\in\mathcal{F}$ such that its OLSE for the full horizon $[T]$ is the solution to (8) assuming BR $b^{1}$. The resulting controls are $(\mathbf{u}^{L1},\mathbf{u}^{F1})$, where $\mathbf{u}^{F1}=b^{\star}(\mathbf{u}^{L1})$. We construct $b^{1}$ such that
$\mathbf{u}^{L1}=\mathbf{v}^{L},\quad\mathbf{u}^{F1}=\mathbf{v}^{F}.$ (21)
Then, by construction, there exists such a $b^{1}$ for which the control played $u^{L1}_{0:\tau-1}$ achieves a lower cost than $u^{L\star}_{0:\tau-1}$ when evaluated over the horizon $0,\tau)$, directly implying (20). ∎
Proposition 1 shows that for the interval $0,\tau)$ there exists another belief $b^{1}\in\mathcal{F}$ that yields a lower leader cost than the true BR $b^{\star}$. Furthermore, since $b^{\star}$ does not minimize (15), multiple beliefs in $\mathcal{F}$ may outperform it. While the assumption that $b^{\star}$ does not minimize (15) may appear restrictive, it is generally the case that the solution to the full-horizon game $[T]$ need not minimize the cost over a subset of the horizon. Example 1 shows an instance where this is indeed the case.
While Proposition 1 establishes that an alternative BR belief may yield a lower leader cost over $0,\tau)$, we next analyze the interval $\tau,T]$. After the BR update at $\tau$, we compare the leader’s cost when it assumes a belief $b^{2}\in\mathcal{F}$ to when it re-optimizes under the true BR belief $b^{\star}$.
Proposition 2.
Under BR update (12) with update time $\tau~<~T$ and OL information structure (4), if the states $x^{1}_{\tau}=x^{\star}_{\tau}$ as defined in (14). Then,
$J^{L}_{\tau:T}(b^{\star})\leq J^{L}_{\tau:T}(b^{2}),\quad\forall b^{2}\in\mathcal{F}.$ (22)
Proof.
The result follows directly from the optimality condition of the OL solution (18) under the problem (11) setup. Since $\hat{u}^{L\star}_{\tau:T-1}$ minimizes the leader’s cost $J^{L}_{\tau:T}$ under the true BR $b^{\star}$ for the new “initial” state $x^{\star}_{\tau}\in{\mathbb{R}}^{n}$. Then, for any other $\hat{u}^{L}_{\tau:T-1}\in{\mathbb{R}}^{m_{L}(T-\tau)}$ announced at $\tau$ and state $x^{\star}_{\tau}$, the cost will be greater. Since the belief $b^{2}\in\mathcal{F}$ induces a corresponding control sequence $\hat{u}^{L2}_{\tau:T-1}$, which yields (22). ∎
Proposition 2 provides a sufficient condition for the optimality of the true BR belief. In general $x^{1}_{\tau}=x^{\star}_{\tau}$ will not be true when the leader’s BR belief over horizon $0,\tau)$, given by $b^{1}\in\mathcal{F}$, is not the true follower BR $b^{\star}\in\mathcal{F}$. Therefore, Proposition 2 implies that assuming the true follower BR $b^{\star}$ during $\tau,T]$ does not necessarily produce the lowest leader cost over $\tau,T]$.
Combining Propositions 1 and 2, we conclude that the answer to Problem 1 is negative in general. For Stackelberg games with a BR update (12), knowing the true follower BR may not result in the lowest cost over $[T]$. Moreover, updating to the true follower’s BR at $t=\tau$ will not guarantee obtaining a lower cost for the leader. Thus, in general, (17) will not hold for all $(b^{1},b^{2})\in\mathcal{F}$. We demonstrate this claim in an example.
Example 1.
Consider a LQ Stackelberg game (App. -A) with scalar state, scalar leader and follower controls, and OL information structure  (4). The state dynamics is
$x_{t+1}=1.7x_{t}+1.4u^{L}_{t}+0.5u^{F}_{t},\quad\forall t\in[T-1].$ (23)
The time horizon is $T=5$, update time is $\tau=3$, and the initial state is $x_{0}=7.6$. The leader’s cost function (43) has parameters $(Q^{L},R^{L})=(16,17)$ and the follower’s cost function (43) with true BR, $b^{\star}$, is given by $(Q^{F},R^{F})=(7,19)$. Let $b^{\prime}\in\mathcal{F}$ denote the follower’s BR if $(Q^{F},R^{F})=(8,9)$. We assume that $b^{1}=b^{2}=b^{\prime}$.
The leader first calculates the strategies $\mathbf{u}^{L\star}$ and $\mathbf{u}^{L^{\prime}}$ under beliefs $b^{\star}$ and $b^{\prime}$ (App. -B). At update time $\tau$, the leader’s cost under both BR beliefs is
$J^{L}_{0:2}(b^{\star})~=~1443.18,\quad J^{L}_{0:2}(b^{\prime})~=~1227.72$ (24)
Thus, $J^{L}_{0:2}(b^{\prime})<J^{L}_{0:2}(b^{\star})$, which verifies Proposition 1’s claim.
Then, at $t=\tau$ the state under BR beliefs $b^{\star}$ is $x^{\star}_{3}~=~1.22$ and the state under BR belief $b^{\prime}$ is $x^{\prime}_{3}~=~2.13$. The updated leader controls are $\hat{u}^{L\star}_{3:4}$ and $\hat{u}^{L^{\prime}}_{3:4}$ (App. -B). We verify that the OLSE of this problem is time-inconsistent since $u^{L\star}_{3:4}~\neq~\hat{u}^{L\star}_{3:4}$.
Finally, the cost for the horizon $[3,5]$ under belief $b^{\star}$ and $b^{\prime}$ is
$J^{L}_{3:5}(b^{\star})=50.72,\quad J^{L}_{3:5}(b^{\prime})=162.88$ (25)
Despite the cost for the horizon $[3,5]$ being greater for $b^{\prime}$, this belief still obtains lower total cost over $[T]$ since $J^{L}_{0:5}(b^{\star})~=~1493.9$ and $J^{L}_{0:5}(b^{\prime})~=~1390.6$.
IV-B Feedback Stackelberg equilibrium (FSE) under BR updates
This section considers whether FSE under follower BR updates satisfies (17). We first adapt the FSE optimality condition [6, Sec.3.6] under BR updates. When the leader solves the Stackelberg game (8) against a follower whose BR is $b^{\star}\in\mathcal{F}$, the leader selects the control FB gain $K^{L\star}_{t}$ that satisfies
$K^{L\star}_{t}\in\mathop{\rm argmin}_{K^{L}_{t}\in\mathcal{K}^{L}_{t}}J^{L}_{t:T}(x_{t},K^{L}_{t},K^{L\star}_{t+1:T-1},\\ b^{\star}(K^{L}_{t}),K^{F\star}_{t+1:T-1}),\quad\forall t\in[T],$ (26)
and $K^{F\star}_{t}=b^{\star}\left(K^{L\star}_{t}\right)$ is the follower’s BR (6). The optimality condition (26) implies that the FB strategy at time $t$ is optimal against leader strategies where only $K^{L}_{t}$ can be changed and $K^{L}_{\hat{t}}$ is fixed for time steps $\hat{t}>t$. This optimality definition ensures that unlike the OLSE, FSE is time-consistent [11, 23]. However, we will show that this is still not enough to affirmatively answer Problem 1.
We now define another property of the FSE that will be relevant in the discussion of whether assuming the true BR will lead to obtaining the lowest cost.
Definition 3 (Markov Perfect FSE (MPFSE) [25]).
A joint FB strategy $\big(\mathbf{K}^{L\star},\mathbf{K}^{F\star}\big)$ is a MPFSE if for all $t\in[T-1]$, in addition to satisfying (26), the leader’s strategy $K^{L\star}_{t:T-1}$ satisfies
$J^{L}_{t:T}\Big(x_{t:T},K^{L\star}_{t:T-1},b^{\star}(K^{L\star}_{t:T-1})\Big)\\ \leq J^{L}_{t:T}\Big(x_{t:T},K^{L}_{t:T-1},b^{\star}(K^{L}_{t:T-1})\Big),\ \forall K^{L}_{t:T-1}\in\mathcal{K}^{L}_{t:T-1},$ (27)
and $K^{F\star}_{t}=b^{\star}\left(K^{L}_{t:T-1}\right)$ is the follower’s BR (6).
MPFSE is a stronger condition than FSE, which only implies time-consistency [23]. In addition to ensuring that the the joint strategy is consistent when resolved in future time steps, MPFSE also ensures that $K^{L\star}_{t:T-1}$ is optimal in the set of all leader FB strategies from time $t$ to T1T-1 [25, Chp.13.2].
Markov perfect Nash equilibrium exists for certain subsets of LQ dynamic games [26, 27], while MPFSE’s existence is less well explored. In [28], MPFSE is shown to exist for two player zero-sum Stackelberg games.
Since the FSE is time-consistent, when the belief over the horizon remains constant, $b^{1}=b^{2}=b$, the sequence of gains $\hat{\mathbf{K}}^{L}$(12) will be the same as $\mathbf{K}^{L}$ which solve (26) under the belief $b\in\mathcal{F}$.
Proposition 3.
Consider the BR update (12) with FB information structure (4). If the FSE is not Markov perfect, there exist a set of beliefs $(b^{1},b^{2})\in\mathcal{F}$ that yield a lower cost than the true belief $b^{\star}\in\mathcal{F}$,
$J^{L}_{0:\tau-1}(b^{1})+J^{L}_{{\tau:T}}(b^{2})<J^{L}_{0:\tau-1}(b^{\star})+J^{L}_{{\tau:T}}(b^{\star}).$ (28)
Proof.
If the FSE is not Markov perfect, this implies that there exists some sequence of FB gains $\mathbf{K}^{L^{\prime}}\neq\mathbf{K}^{L\star}$ such that they minimize $J^{L}_{0:T}$ (5) over all other possible gain sequences with follower BR $K^{F^{\prime}}_{t}=b^{\star}(K^{L^{\prime}}_{t})$. We select some $b^{1}\in\mathcal{F}$ such that the solution to (8) under belief $b^{1}$ is denoted $\left(\mathbf{K}^{L1},\mathbf{K}^{F1}\right)$ and $K^{F1}_{t}=b^{\star}(K^{F1}_{t})$. We assume that, in addition, this $b^{1}$ satisfies
$K^{L1}_{0:\tau-1}=K^{L^{\prime}}_{0:\tau-1}\text{ and }K^{F1}_{0:\tau-1}=K^{F^{\prime}}_{0:\tau-1}.$ (29)
Similarly, we can define another $b^{2}\in\mathcal{F}$ such that the solution to (11) under belief $b^{2}$ is $\left(\hat{K}^{L2}_{\tau:T-1},\hat{K}^{F2}_{\tau:T-1}\right)$ and $\hat{K}^{F2}_{t}=b^{\star}(\hat{K}^{F2}_{t})$. In addition, this $b^{2}$ satisfies
$\hat{K}^{L2}_{\tau:T-1}=K^{L^{\prime}}_{\tau:T-1}\text{ and }\hat{K}^{F2}_{\tau:T-1}=K^{F^{\prime}}_{\tau:T-1}.$ (30)
Thus, we have constructed a belief update $(b^{1},b^{2})\in\mathcal{F}^{2}$ that obtains lower cost than the true belief $b^{\star}$, implying (28). ∎
Proposition 3 implies that some BR beliefs other than the true BR may yield lower leader lost. This is because the FSE is generally not Markov perfect. Since the true BR FSE does not minimize the $[T]$ horizon cost over all possible gain sequences, there may be multiple pairs $(b^{1},b^{2})\in\mathcal{F}^{2}$ that obtain lower cost than $b^{\star}$. The claim made in this proposition is illustrated in a numerical example.
Example 2.
This example presents a LQ Stackelberg game (App. -A) with FB information structure (4) where assuming some $b^{\prime}\in\mathcal{F}$ achieves a lower cost than $b^{\star}\in\mathcal{F}$. Consider the scalar game dynamics
$x_{t+1}=1.4x_{t}+1.7u^{L}_{t}+1.7u^{F}_{t},\quad\forall t\in[T-1],$ (31)
with a horizon $T=8$, update time $\tau=3$, and initial condition $x_{0}=-5.6$. The leader’s cost (43) is defined by the matrices $(Q^{L},R^{L})=(7,16)$. The follower’s true BR $b^{\star}$ is given by $(Q^{F},R^{F})=(4,24)$. We consider the case where the leader also assumes a BR $b^{1}=b^{2}=b^{\prime}$ defined by $(Q^{F},R^{F})=(29,12)$. Because the FSE is time-consistent and the beliefs remain constant, the sequence of leader gains under beliefs $b^{\star}$ and $b^{\prime}$, respectively, are $\mathbf{K}^{L\star}$ and $\mathbf{K}^{L^{\prime}}$ and the follower’s BR in either case is $\mathbf{K}^{F\star}$ and $\mathbf{K}^{F^{\prime}}$. All numerical values are provided in App. -C. After playing the different strategies, the leader observes costs
$J^{L}_{0:7}\left(b^{\star}\right)=347.2,\quad J^{L}_{0:7}\left(b^{\prime}\right)=299.4$ (32)
Which shows that the case where the leader assumed the wrong follower BR obtained a lower cost than the true BR.
V Simulation results
In this section, we provide a more physically meaningful example of how the propositions presented in Section IV apply in a conflict-avoidance scenario.
V-A Simulation set up
The framework given below can be utilized to model a trajectory game between two agents where limited communication is available, such as a human-robot encounter. We will assume that each player is a two dimensional double integrator with dynamics
$x^{i}_{t+1}=\begin{bmatrix}1&1\\ 0&1\end{bmatrix}x^{i}_{t}+\begin{bmatrix}0.5\\ 1\end{bmatrix}u^{i}_{t},\ \forall t\in[T-1],\ \forall i\in\{L,F\}.$ (33)
The game state is made up of the leader and follower’s individual states, as well as the players’ reference trajectories, $r^{i}_{t}$; $x_{t}=\left[{x^{L}_{t}}^{\top},{x^{F}_{t}}^{\top},{r^{L}_{t}}^{\top},{r^{F}_{t}}^{\top}\right]^{\top}$. These reference trajectories evolve according to the dynamics
$r^{i}_{t+1}=\Sigma^{i}r^{i}_{t}\quad\forall t\in[T-1],\ \forall i\in\{L,F\}.$ (34)
The leader’s objective throughout the game is to simultaneously avoid the follower while tracking its reference trajectory. Thus, its state cost matrix is the $4\times 4$ block matrix
$Q^{L}=\begin{bmatrix}Q^{L1}&Q^{L2}&Q^{L3}&0\\ {Q^{L2}}^{\top}&\varepsilon I_{2\times 2}&0&0\\ {Q^{L3}}^{\top}&0&Q^{L4}&0\\ 0&0&0&0\end{bmatrix}\in{\mathbb{R}}^{8x8},$ (35)
where $Q^{L2}\succ 0$, $Q^{L3}\prec 0$, $\varepsilon>0$ and the remaining matrices are positive definite (PD) and set to ensure $Q^{L}\succeq 0$. The control cost matrix is $R^{L}\succ 0$.
We consider three possible intentions for the follower: either is is tracking the leader (T), it is indifferent to it (I), or it is avoiding the leader (A). In all cases the control cost matrix is $R^{F}\succ 0$. The state cost matrices in each case are
$Q^{F}{(T)}=\begin{bmatrix}Q^{F1}{(T)}&Q^{F3}{(T)}&0&0\\ {Q^{F3}{(T)}}^{\top}&Q^{F2}{(T)}&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}\in{\mathbb{R}}^{8x8},$ (36)
where $Q^{F3}{(T)}\prec 0$ and the remaining matrices are PD and set to ensure $Q^{F}{(T)}\succeq 0$. When the follower is indifferent:
$Q^{F}{(I)}=\begin{bmatrix}0&0&0&0\\ 0&Q^{F2}{(I)}&0&Q^{F4}{(I)}\\ 0&0&0&0\\ 0&{Q^{F4}{(I)}}^{\top}&0&Q^{F5}{(I)}\end{bmatrix}\in{\mathbb{R}}^{8x8}.$ (37)
In this case $Q^{F4}{(I)}\prec 0$ and remaining matrices are PD and ensure $Q^{F}{(I)}\succeq 0$. Finally, when the follower is avoiding the leader its state cost is given by:
$Q^{F}{(A)}=\begin{bmatrix}\alpha I_{2\times 2}&Q^{F3}{(A)}&0&0\\ {Q^{F3}{(A)}}^{\top}&Q^{F2}{(A)}&0&Q^{F4}{(A)}\\ 0&0&0&0\\ 0&{Q^{F4}{(A)}}^{\top}&0&Q^{F5}{(A)}\end{bmatrix}\in{\mathbb{R}}^{8x8}.$ (38)
Here, $Q^{F4}{(A)}\prec 0$, $\alpha>0$ and all other matrices are PD and ensure $Q^{F}{(A)}\succeq 0$.
Intention estimation algorithm.We compare the leader’s performance under four control strategies: assuming one of three fixed follower intentions (T,I,A) or using an adaptive control (Ad). In the adaptive control, the leader updates its assumption about the follower’s BR every $\tau$ steps using a Multiple Model Adaptive Estimator (MMAE) [29]. At time $t\in[T]$, the BR belief is
$b_{t}=\mathop{\rm argmax}_{b}P\big(b|x_{t}\big),$ (39)
where $P\big(b|x_{t}\big)$ is the probability of BR $b\in\mathcal{F}$ based on the current state observation $x_{t}\in{\mathbb{R}}^{n}$. The probabilities are initialized to be equal, $P\big(b|x_{0}\big)=\frac{1}{3}$. Then, they are updated according to a Bayesian update rule
$P\big(b|x_{t}\big)=\frac{P\big(x_{t}|b\big)\cdot P(b)}{\sum_{\beta=1}^{3}P\big(x_{t}|\beta\big)\cdot P(\beta)},$ (40)
where $P(b)=P\big(b|x_{t-1}\big)$. The model probability for each assumption $b$ is determined by applying a softmin function to the corresponding residual norm
$P\big(x_{t}|b\big)=\frac{\exp(-\|e_{t}({b})\|_{2})}{\sum_{\beta=1}^{3}\exp(-\|e_{t}({\beta})\|_{2})}$ (41)
Where the residual is defined as $e_{t}(b)=x_{t}-\hat{x}_{t}(b)$, with $\hat{x}_{t}(b)$ being the propagated dynamics (1) if the follower had BR $b$.
Monte Carlo simulation setup For each experiment we run a total of 1000 Monte Carlo simulations for a game horizon $T=20$. In each simulation, the initial state $x_{0}$ is sampled from a uniform distribution such that $\|x_{0}\|_{\infty}\leq 20$ and the reference trajectories are defined by $\Sigma^{i}=\sigma I^{2\times 2}$, where $0<\sigma\leq 1$ is sampled uniformly.
V-B OL information structure
V-B1 Optimality against other beliefs
This case study verifies that an incorrect BR belief may lead to obtaining the lowest cost. The following matrix shows for each follower ground truth, what percentage of times each BR belief obtains the lowest cost.
Figure 2: Percent of simulations with lowest cost achieved by each BR belief under OL information structure when $\tau=1$
V-B2 Update time case study
We examine how the update time $\tau$ affects how often each BR belief yields the lowest cost. Figure 3 shows this for the case when the follower’s true BR is avoiding.
Figure 3: Percentage of simulations where each BR belief obtains the lowest cost for $\tau=1,2,5,10,20$ under OL information structure
V-C Feedback information structure
V-C1 Optimality against other beliefs
This case study is identical to Figure 2 in setup, but the controls for both players come from the FSE solution (App. -A).
Figure 4: Percent of simulations with lowest cost achieved by BR beliefs under FB information structure and $\tau=1$
V-C2 Estimation horizon case study
We examine how the update time $\tau$ affects how often each BR belief yields the lowest cost. Figure 5 shows the case when the follower’s true BR is avoiding.
Figure 5: Percentage of simulations where each BR belief obtains the lowest cost for $\tau=1,2,5,10,20$ under FB information structure
V-D Cost difference
[…]
% $\%\text{ higher cost}=\frac{J^{L}_{0:T}(\hat{b})-\min_{b}J^{L}_{0:T}(b)}{\min_{b}J^{L}_{0:T}(b)}\cdot 100$ (42)
Figure 6: Average percentage by which each belief’s cost exceeds the best model’s cost for each true BR and $\tau=1$ under OL and FB information structures
V-E Discussion
This section discusses the results presented in Sections V-B and V-C. The first key observation is that Figures 2 and 4 show that the lowest cost is not always achieved by the true BR belief. Under both information structures, when the follower is either tracking or avoiding the leader, some alternative beliefs occasionally yield better performance.
Another important insight from Figures 2 and 4 is that the performance associated with assuming a particular belief depends on the set of beliefs under consideration. For example, when the follower is indifferent, assuming the true BR consistently results in the lowest cost. However, expanding the set of possible behaviors may alter this outcome. This aspect is a subject for future investigation.
Figures 3 and 5 explore the relationship between the update interval $\tau$ and the frequency with which the true BR belief is outperformed by others. In the OL case, we observe that more frequent updates lead to more instances in which alternative models perform better, with $\tau=T=20$ corresponding to the optimality condition of the OLSE (18). In contrast, in the FB case, the number of times each model achieves the lowest cost remains constant across update intervals. This behavior aligns with the time-consistency property of the FSE: since each recomputation produces the same solution, the relative performance of the strategies remains unaffected by $\tau$.
Finally, Figure 6 shows that, across all simulations, the true BR belief remains closest to the minimum cost overall, while the cost gap with incorrect BR beliefs is, on average, quite large. This finding is noteworthy: although some beliefs occasionally outperform the true BR, this does not translate into comparable overall cost levels. In all cases, the adaptive solution also performs near-optimally on average.
VI Conclusion
This paper examines the optimality conditions of Stackelberg OL and FB equilibria when the leader’s belief about the follower’s BR is updated over the game horizon. We prove sufficient conditions for the optimality of the true BR belief under the OL information structure and discuss its connection to the time-inconsistency of the OLSE. Under FB information structure, the SE is not necessarily Markov perfect, implying that when the leader holds incorrect BR beliefs, the resulting cost may, in some cases, be lower. Finally, we illustrate how these inconsistencies arise in a LQ game modeling a conflict-avoidance scenario. Future work will focus on characterizing the BR beliefs that may yield a lower cost than the true BR and identifying the conditions under which such outcomes occur.